### v2

@@ 4 Engineering
We emphasize several critical points that must be carefully implemented in order to obtain an efficient code. In both geometric heuristics and local search, the majority of computational time is spent determining whether two items overlap.
@@ 4.1 Fast Crossing Detection
The data structure also includes a decomposition of the item boundary into monotone polylines, going from left to right (plus a boolean value telling whether the item is above or below the polyline). We call them chains. Each chain is also assigned its own bounding box. Convex chains have only two chains. Intersections between chains are computed in linear time. If the chains of two items cross, then the items overlap. In cases where the chains share common points without crossing (such as a shared vertex) special attention is required. These degenerate cases are particularly significant, as they correspond to configurations where the items are positioned to touch but not overlap, which is often the desired outcome.
@@ 4.3 Push Routine and Issues
The push routine is designed to allow items to slide along the boundaries of other packed items and the container. To achieve this, we consider multiple integer vectors $v$ such that the dot product $u\cdot v$ is strictly positive, then translate the item as far as possible along the ray passing through $p$ in direction $v$. If a new valid position is found, we update the position $p$ and test another direction $v$. After testing a few directions $v$ without successfully moving the item $i$, we conclude the push routine.
We first choose the integer vector $u$, for example a normal vector to the diameter of the item. Let $u^{\prime}$ be a vector perpendicular to $u$. We try vectors $v$ of the form $v=u+\alpha u^{\prime}$ for integer $\alpha$ from $-8$ to $8$.
@@
The translation in the direction $v$ starts by computing the smallest factor $2^{k}$ for integer $k$, such that $p+2^{k}v$ lies outside the container. We then repeat the following procedure. We decrement $k$ until we find a position $p+2^{k}v$ where the item $i$ can be placed or $k$ becomes 00. If such a position is found, we update $p$ to the new position $p+2^{k}v$. We then repeat the procedure for the same vector $v$ and smaller values of $k$ until $p+v$ is not a valid position.
A problem that may arise is that, since $v$ is an integer vector, its length may be large, potentially leaving empty space between $p$ and $p+v$. To address this, we continue the process by investigating new positions along $p+2^{k}v$ for progressively smaller values of $k$ (e.g., $k=-1$, $k=-2$, and so on). Since only integer translations are allowed, we may need to round the coordinates of $2^{k}v$ when $k$ is negative. This introduces two potentially significant issues that need to be addressed. Both issue arise from the fact that division and rounding can alter the direction of the vector.
The first issue happens because $\textrm{round}(2^{k}v)$ is sometimes valid for much higher values of $k$, which takes a very long time. To solve this, we set $v$ to $\textrm{round}(2^{k}v)$. As the direction is slightly altered, we recompute the smallest integer $k$, such that $p+2^{k}v$ lies outside the container and start decreasing it as before.
The second issue is that $u\cdot\textrm{round}(2^{k}v)$ may become negative. If that happens, we stop the calculation and test other different directions $v$. Despite its apparent complexity, the push routine we implemented, which proceeds by discrete jumps and subsequently only calls the overlapping test, is surprisingly efficient.

### v3

@@ 4 Engineering
We emphasize several critical points that must be carefully implemented in order to obtain efficient code. In both geometric heuristics and local search, the majority of computational time is spent determining whether two items overlap.
@@ 4.1 Fast Crossing Detection
The data structure also includes a decomposition of the item boundary into monotone polylines, going from left to right (plus a boolean value telling whether the item is above or below the polyline). We call them chains. Each chain is also assigned its own bounding box. Convex items have only two chains. Intersections between chains are computed in linear time. If the chains of two items cross, then the items overlap. In cases where the chains share common points without crossing (such as a shared vertex) special attention is required. These degenerate cases are particularly significant, as they correspond to configurations where the items are positioned to touch but not overlap, which is often the desired outcome.
@@ 4.3 Push Routine and Issues
The push routine is designed to allow items to slide along the boundaries of other packed items and the container. To achieve this, we consider multiple integer vectors $v$ such that the dot product $u\cdot v$ is strictly positive, then translate the item as far as possible along the ray passing through $p$ in direction $v$. If a new valid position is found, we update the position $p$ and test another direction $v$. After testing a few directions $v$ without successfully moving item $i$, we end the push routine.
We first choose the integer vector $u$, for example a normal vector to the diameter of the item. Let $u^{\prime}$ be a vector perpendicular to $u$. We try vectors $v$ of the form $v=u+\alpha u^{\prime}$ for integer $\alpha$ from 8-8 to $8$.
@@
The translation in the direction $v$ starts by computing the smallest factor $2^{k}$ for integer $k$, such that $p+2^{k}v$ lies outside the container. We then repeat the following procedure. We decrement $k$ until we find a position $p+2^{k}v$ where item $i$ can be placed or $k$ becomes 0. If such a position is found, we update $p$ to the new position $p+2^{k}v$. We then repeat the procedure for the same vector $v$ and smaller values of $k$ until $p+v$ is not a valid position.
A problem that may arise is that, since $v$ is an integer vector, its length may be large, potentially leaving empty space between $p$ and $p+v$. To address this, we continue the process by investigating new positions along $p+2^{k}v$ for progressively smaller values of $k$ (e.g., k=1k=-1, k=2k=-2, and so on). Since only integer translations are allowed, we may need to round the coordinates of $2^{k}v$ when $k$ is negative. This introduces two potentially significant issues that need to be addressed. Both issues arise from the fact that division and rounding can alter the direction of the vector.
The first issue happens in the following case. The vector $\textrm{round}(2^{k}v)$ is valid but since it is the first time that we test this rounded direction, it might happen that the item can be translated by $2^{r}\textrm{round}(2^{k}v)$ with a large integer $r$. To precise the issue, let us consider for instance an example with $v=(2,3)$. We assume that the item cannot be translated by $v$. Then we round it to $2^{-1}v$ which becomes $(1,1)$. It is possible that in this new direction we can translate the item from $\alpha v$ with $\alpha$ going from $1$ to, for instance, $10^{8}$. If we repeat translations of vector $(1,1)$ until reaching an obstacle, it will take a very long running time. To avoid this problem, each time that the direction is altered by a rounding, we proceed as if it was a new direction: we set $v$ to $\textrm{round}(2^{k}v)$ and we recompute the smallest integer $k$, such that $p+2^{k}v$ lies outside the container and start decreasing it as before.
The second issue is that $u\cdot\textrm{round}(2^{k}v)$ may become negative. If that happens, we stop the calculation and test other different directions $v$. Despite its apparent complexity, the push routine we implemented, which proceeds by discrete jumps and subsequently calls only the overlapping test, is surprisingly efficient.